OpenStudy (anonymous):

Which of the following is the radical expression of 2 times d to the seven tenths power?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[2\sqrt[10]{d^7}\]

OpenStudy (anonymous):

\[\Large a \cdot {x^{n/m}} = a \cdot \sqrt[m]{{{x^n}}}\]